# Matryoshka: Learning to Drive Black-Box LLMs with LLMs

### v1

@@ Appendix B Additional Related Works
Appendix C Dataset and Task Details
C.1 LaMP: Personalization
@@
We employ the Language Model Personalization (LaMP) benchmark (Salemi et al., 2023), an open-source benchmark specifically designed to train and evaluate the capability of language models in generating personalized content. LaMP encompasses a diverse set of tasks (with LaMP-2 comprising two tasks, LaMP-2N, and LaMP-2M), covering both personalized text classification and generation tasks. The dataset statistics are presented in Table 5 for a clear overview of its structure. Below are detailed descriptions of each task:
@@
Table 5: Dataset statistics of five different personalization tasks (LaMP-1, 2N, 2M, 3, and 4) from the LaMP benchmark (Salemi et al., 2023).
@@
LaMP-1 Classification 9682 2500 2500 $51.40\pm 5.72$ $90.61\pm 53.87$ 2
LaMP-2N Classification 5914 1052 1274 $65.40\pm 12.29$ $306.42\pm 286.65$ 15
LaMP-2M Classification 5073 1410 1557 $92.39\pm 21.95$ $86.76\pm 189.52$ 15
LaMP-3 Classification 20000 2500 2500 $145.14\pm 157.96$ $188.10\pm 129.42$ 5
LaMP-4 Generation 12527 1925 2376 $30.53\pm 12.67$ $9.78\pm 3.10$ $287.16\pm 360.62$
C.2 Reasoning: GSM8K
GSM8K (Cobbe et al., 2021) is a dataset focused on high school-level mathematical reasoning. The numerical reasoning tasks within this dataset typically consist of a descriptive scenario followed by a culminating question. Answering these questions requires performing multi-step mathematical calculations based on the context provided in the description.
C.3 Planning: ALFWorld
AlfWorld (Shridhar et al., 2020) is a comprehensive suite of synthetic, text-based environments set within a virtual household, featuring six distinct task types: Pick, Clean, Heat, Cool, Examine, and Pick Two. Each task presents a unique high-level objective (e.g., “put a vase in the safe”) that requires the agent to navigate and interact with various objects or receptacles (e.g., go to shelf 6, clean apple). To accomplish the assigned task, the agent must execute a series of actions to achieve the specified goal. However, the challenge lies in the object’s potential location - it could be in any of over 50 possible places within a given task instance - necessitating sequential exploration of each location by the agent. Consequently, the complete action sequence may encompass more than 50 discrete actions, posing a considerable challenge to the agent’s capabilities and efficiency.
Appendix D Baseline Details
D.1 LaMP: Personalization
We compare our proposed Matryoshka with several competitive baselines, encompassing both one-stage and two-stage methods. For all baseline approaches, we employ a consistent prompt template and utilize BM25 as the default retrieval mechanism across all experiments.
@@ D.1 LaMP: Personalization
For our ablation study, we primarily compare Matryoshka with the following ablated baseline:
  • Matryoshka w/o IGO utilizes the controller model Llama-3-8B-Instruct to first generate a summary of the user’s retrieved history data. It then combines this summary with the input question as prompts for the environment model gpt-4o-mini to generate the final answer.
D.2 GSM: Reasoning
@@
  • Matryoshka w/o IGO first utilizes a vanilla LLaMA3-8B-Instruct to break down the problem into sub-questions, and then gpt-3.5-turbo provide solutions based on both the main problem and the decomposed sub-questions.
@@
D.3 Alfworld: Planning
We compare Matryoshka with several strong baselines in the planning task, encompassing both one-stage and two-stage approaches. For all baselines, we employ gpt-3.5-turbo as the black-box model for task execution. The ablated baselines (two-stage) include w/o Guidance Optimization, w/o $1^{\text{st}},2^{\text{nd}}$-round IGO, and w/o $2^{\text{nd}}$-round IGO. Additional baselines (one-stage) include BUTLER (Shridhar et al., 2020), ReAct (Yao et al., 2023), Reflexion (Shinn et al., 2023), and AdaPlanner (Sun et al., 2024a).
@@ D.3 Alfworld: Planning
    • w/o $1^{\text{st}},2^{\text{nd}}$-round IGO indicates a LLaMA3-8B-Instruct model that has undergone supervised fine-tuning on a limited amount of training data.
    • w/o $2^{\text{nd}}$-round IGO denotes the LLaMA3-8B-Instruct model further trained using DPO on {positive, negative} pairs from the training set, building upon the supervised fine-tuned model.
  • BUTLER (Shridhar et al., 2020) is an agent that initially learns to perform abstract tasks in TextWorld through Imitation Learning (IL) and subsequently transfers the acquired policies to embodied tasks in ALFWorld.
  • ReAct (Yao et al., 2023) is a general paradigm that combines reasoning and acting with language models to solve diverse language reasoning and decision-making tasks.
  • Reflexion (Shinn et al., 2023) employs verbal reinforcement to help agents learn from prior failures.
  • AdaPlanner (Sun et al., 2024a) is a closed-loop planning method where the LLM plays two roles, planner and refiner. It leverages code-based prompting for precise planning and refinement.
Appendix E Implementation Details
E.1 Hardware and Software
@@
E.2 LaMP: Personalization
E.2.1 Algorithm Details
We formalize the personalization problem within the context of our proposed Matryoshka framework. Specifically, we employ the controller model Llama-3-8B-Instruct to analyze the user’s retrieved history data and generate an informative and clear intermediate summary. This summary is then combined with the input question as prompts for the environment model gpt-4o-mini to derive the final answer. To enhance control capabilities, we utilize online DPO to optimize the controller model Llama-3-8B-Instruct.
@@ E.2.1 Algorithm Details
To prevent overfitting and reward hacking, the interaction stage processes the entire training dataset once for all personalization tasks. We sample at most two contrastive pairs for each training data point. In cases where no positive generation exists for some data points, we utilize a more powerful model, such as gpt-4o, to produce several strong intermediate generations, thereby increasing the likelihood of obtaining positive samples. We employ LoRA (Low-Rank Adaptation), a parameter-efficient method, to update the controller model Llama-3-8B-Instruct. LoRA is well-suited for personalization tasks, allowing efficient and effective optimization of the controller model. We utilize DPO for optimization.
E.2.2 Hyperparameter Configurations
We set the maximum sequence length for generated solutions to 512 tokens across all tasks and scenarios. The controller model is Llama-3-8B-Instruct, while the environment model is gpt-4o-mini for the primary tasks and gpt-3.5-turbo for specific ablation studies. For each user, we retrieve the minimum of $k$ and the total number of user profiles as historical profiles. The value of $k$ varies by dataset: all profiles for LaMP-1, 120 for LaMP-2N, 150 for LaMP-2M, 30 for LaMP-3, and 50 for LaMP-4. These retrieved profiles are utilized in generating intermediate solutions. Comprehensive prompt templates and additional details are provided in Appendix G.
To prevent overfitting and reward hacking, we iterate through the entire training dataset only once. For each data point, we perform ten interactions, generating ten distinct intermediate solutions with a temperature setting of 1.0. If no positive samples are identified, we employ the more advanced gpt-4o to generate an additional five intermediate solutions, applying the same criteria to evaluate their positivity. Consequently, each data point results in at least ten intermediate generations. To further mitigate overfitting and reward hacking, we sample a maximum of two contrastive pairs per data point. The total number of contrastive pairs sampled during the interaction stage is as follows: 5410 for LaMP 1, 2850 for LaMP-2M, 2548 for LaMP-2N, 4320 for LaMP-3, and 12518 for LaMP-4, respectively.
@@ E.2.2 Hyperparameter Configurations
E.3 GSM8K: Reasoning
Following the PAL framework (Gao et al., 2023), we employ code-style LLM prompts to facilitate the conversion of mathematical problems into executable code, thereby augmenting the model’s problem-solving capabilities. Unlike PAL, which directly translates mathematical problems into code, Matryoshka first assists GPT in decomposing the problem into more manageable sub-problems. This decomposition allows GPT to more effectively convert these simpler sub-problems into code, enhancing both the correctness and stability of the generated code. Additionally, since Matryoshka is responsible solely for high-level planning without engaging in low-level execution, we can train on the GSM8K dataset and evaluate on the GSM-Hard dataset. Both datasets comprise similar problem types, with GSM-Hard featuring more intricate numerical calculations.

### v2

@@ B.4 M-Pilot with other Iterative Training Method
M-Pilot (IDPO) 0.931 0.964 0.761 0.801
M-Pilot (ISimPo) 0.908 0.950 0.731 0.789
Table 9: Additional experimental results across six planning tasks from AlfWorld.
BUTLER [39] 46.00 39.00 74.00 100.00 22.00 24.00 37.00
ReAct [55] 37.50 64.52 69.57 42.86 38.89 17.65 47.76
Reflexion [38] 50.00 41.94 65.22 52.38 66.67 47.06 52.99
AdaPlanner [41] 100.00 93.55 78.26 95.24 66.67 88.24 88.06
M-Pilot (IDPO) 100.00 93.55 100.00 95.24 100.00 88.24 96.27
M-Pilot (ISimPo) 100.00 93.55 95.65 95.24 77.78 88.24 92.54
The results presented in Table 7, Table 8, and Table 9, despite being obtained with sub-optimal hyperparameters, still demonstrate that iterative guidance optimization can be seamlessly incorporated into various reinforcement learning methods. Moreover, it consistently outperforms other state-of-the-art baselines, further underscoring the effectiveness and versatility of the framework.
B.5 M-Pilot with other controller model
Table 10: M-Pilot compared with baselines on LaMP benchmark. The Controller model utilizes Qwen2.5-7B-Instruct as the base model, other settings remain consistent with main experiment.
Dataset ($\rightarrow$) LaMP-1 LaMP-2N LaMP-2M LaMP-3 LaMP-4
gpt-4o-mini 0.514 0.513 0.655 0.473 0.413 0.325 0.371 0.673 0.132 0.116 0.992
RAG (k=1) [34] 0.626 0.624 0.733 0.539 0.444 0.378 0.311 0.631 0.141 0.126 1.296
RAG (k=4) [34] 0.632 0.632 0.792 0.611 0.502 0.430 0.272 0.579 0.161 0.146 2.953
PAG [32] 0.624 0.624 0.775 0.559 0.496 0.443 0.316 0.645 0.143 0.130 1.968
M-Pilot (Qwen) 0.640 0.640 0.808 0.579 0.509 0.443 0.301 0.619 0.167 0.153 4.466
w/o IDPO 0.608 0.608 0.777 0.573 0.499 0.432 0.313 0.637 0.156 0.139 2.296
Table 11: M-Pilot compared with baselines on GSM8K dataset using Qwen2.5-7B-Instruct as the controller model.
Dataset ($\rightarrow$) GSM8K GSM-HARD
Method ($\downarrow$) gpt-3.5 4o-mini gpt-3.5 4o-mini
CoT 0.809 0.932 0.406 0.500
Least-to-Most 0.811 0.908 0.425 0.498
PAL 0.802 0.920 0.638 0.748
[…]
M-Pilot (Qwen) 0.936 0.964 0.773 0.800
w/o IDPO 0.932 0.961 0.767 0.799
Table 12: M-Pilot compared with baselines on Alfworld dataset using Qwen2.5-7B-Instruct as the controller model.
Methods ($\downarrow$) Tasks ($\rightarrow$) Pick Clean Heat Cool Exam Pick2 All
BUTLER [39] 46.00 39.00 74.00 100.00 22.00 24.00 37.00
ReAct [55] 37.50 64.52 69.57 42.86 38.89 17.65 47.76
Reflexion [38] 50.00 41.94 65.22 52.38 66.67 47.06 52.99
AdaPlanner [41] 100.00 93.55 78.26 95.24 66.67 88.24 88.06
M-Pilot (Qwen) 100.00 93.55 100.00 80.95 55.56 94.12 88.81
w/o IDPO 100.00 90.32 82.61 80.95 44.44 94.12 83.58
Moreover, from a modular perspective, each component in the pipeline can be flexibly substituted without disrupting the overall framework. In this section, we replace the controller model from LLaMA3-8B-Instruct with Qwen2.5-7B-Instruct and replicate experiments across all domains as presented in the main paper. As shown in Tables 1011, and 12, M-Pilot continues to deliver strong results, achieving up to 1.6%, 7.2%, and 0.8% improvements over the second-best baseline across the respective benchmarks. These results further underscore its broad effectiveness and robustness across different controller models.
B.6 Comparison with Black-Box LLMs Controllers
Table 13: M-Pilot compared to Black-Box LLM Controllers on AlfWorld. For tasks in ALFWorld, we adhered to the setup used in Adaplanner [41].
gpt-3.5 + gpt-3.5 100.00 41.94 100.00 76.19 88.89 88.24 79.85
gpt-4o-mini + gpt-3.5 95.83 45.16 56.52 52.38 5.56 88.24 57.46
M-Pilot + gpt-3.5 100.00 93.55 100.00 95.24 100.00 88.24 96.27
Table 14: M-Pilot compared to Black-Box LLM Controllers on GSM8K dataset.
Method ($\downarrow$) Dataset ($\rightarrow$) GSM8K GSM-HARD
gpt-3.5 + gpt-3.5 0.896 0.734
gpt-4o-mini + gpt-4o-mini 0.948 0.791
M-Pilot + gpt-4o-mini 0.964 0.801
M-Pilot + gpt-3.5 0.931 0.761
Table 15: M-Pilot compared to Black-Box LLM Controllers on the LaMP benchmark.
Dataset ($\rightarrow$) LaMP-1 LaMP-2N LaMP-2M LaMP-3 LaMP-4
gpt-3.5 + gpt-3.5 0.590 0.589 0.790 0.594 0.399 0.325 0.357 0.693 0.166 0.150 3.433
gpt-4o-mini + gpt-4o-mini 0.624 0.624 0.775 0.559 0.496 0.443 0.316 0.645 0.143 0.130 1.968
M-Pilot + gpt-4o-mini 0.640 0.639 0.823 0.607 0.527 0.465 0.277 0.581 0.174 0.160 4.298
M-Pilot + gpt-3.5 0.594 0.593 0.798 0.609 0.469 0.412 0.286 0.599 0.176 0.161 4.222
We further compare our controller-generator framework with directly using a black-box LLM to guide another black-box LLM. As shown in Tables  1314, and 15, our framework consistently delivers either comparable or significantly better results across tasks in planning, reasoning, and personalization. We attribute this to Effective Problem Decomposition with Feedback. As highlighted in our abstract, we treat the black-box LLM within our framework as an ”environment” and the white-box LLM as a “controller”. The white-box LLM decomposes the problem and provides it as input to the black-box LLM. The black-box LLM either interacts with the environment or compares its output against ground truth, returning feedback as a supervisory signal. This feedback helps filter high-quality problem decompositions to train the white-box LLM. As a result, the trained white-box LLM generates problem decompositions that more effectively guide the black-box LLM to solve tasks. In contrast, a black-box LLM alone lacks environment feedback and cannot achieve equally effective problem decomposition.
B.7 Reduced Token Usage
Table 16: API cost on AlfWorld.
Method API cost (\$) Task Performance
M-Pilot 0.818 97.8
Adaplanner [41] (w/o controller) 1.151 88.8
For the AlfWorld task, we follow the settings used in Adaplanner [41], allowing the black-box LLM to reflect up to eight times. The process terminates either upon success or when the maximum number of reflections is reached. We report the black-box API cost and task performance with and without the controller in the table below. All experiments are conducted using gpt-3.5-turbo as the black-box LLM. As shown in Table 16, our method reduces API cost by 30% while still achieving a 9% improvement in task performance. This is attributed to the controller model’s well-structured instructions, which enable the black-box LLM to interact with the environment fewer times while attaining a higher success rate.
Table 17: API cost on GSM8K.
Method API cost (\$) Task Performance
M-Pilot 1.740 93.1
Self-Consistency [49] (w/o controller) 4.946 81.3
For the GSM8K task, we compare our method with the self-consistency [49] approach, where the black-box model generates 8 responses per question and selects the final answer via majority voting. As shown in Table 17, compared to self-consistency, our method reduces API cost by 65% while improving performance by over 12% , thanks to the high-quality guidance provided by the controller model.
Appendix C Additional Related Works
Small LMs Drive LLMs Generation. SuperICL [53] incorporates outputs from smaller language models (LMs) as complementary information for input queries, integrating them into the context provided to black-box LLMs. However, these smaller LMs are fixed and can only support classification tasks that rely on label predictions with associated confidence scores. HYDRA [63] is a retrieval-augmented generation framework that trains a BERT-sized reranker to reorder retrieved passages to better cater to user-specific requirements. Nevertheless, these methods apply only discrete optimization on the prompt through reranking and selection of few-shot demonstrations, which limits the potential improvements achievable via prompt engineering.
RLHF. Proximal policy optimization (PPO) [35] is the predominant deep reinforcement learning method used in RLHF, leading to significant successes in models like InstructGPT [28], ChatGPT [1], and Gemini [31]. However, applying PPO requires extensive effort and resources [8, 13, 45], often beyond the scope of open-source capabilities. To simplify implementation and streamline the training process, recent works [2, 14] have proposed direct preference learning algorithms following the DPO framework [30]. These algorithms bypass the reward modeling step and directly optimize carefully designed loss objectives on the preference dataset, hence the term direct preference learning.
[…]
Appendix D Dataset and Task Details
D.1 LaMP: Personalization
@@
We employ the Language Model Personalization (LaMP) benchmark [34], an open-source benchmark specifically designed to train and evaluate the capability of language models in generating personalized content. LaMP encompasses a diverse set of tasks (with LaMP-2 comprising two tasks, LaMP-2N, and LaMP-2M), covering both personalized text classification and generation tasks. The dataset statistics are presented in Table 18 for a clear overview of its structure. Below are detailed descriptions of each task:
@@
Table 18: Dataset statistics of five different personalization tasks (LaMP-1, 2N, 2M, 3, and 4) from the LaMP benchmark [34].
@@
LaMP-1 Classification 9682 2500 2500 $51.40\pm 5.72$ - $90.61\pm 53.87$ 2
LaMP-2N Classification 5914 1052 1274 $65.40\pm 12.29$ - $306.42\pm 286.65$ 15
LaMP-2M Classification 5073 1410 1557 $92.39\pm 21.95$ - $86.76\pm 189.52$ 15
LaMP-3 Classification 20000 2500 2500 $145.14\pm 157.96$ - $188.10\pm 129.42$ 5
LaMP-4 Generation 12527 1925 2376 $30.53\pm 12.67$ $9.78\pm 3.10$ $287.16\pm 360.62$ -
D.2 Reasoning: GSM8K
GSM8K [10] is a dataset focused on high school-level mathematical reasoning. The numerical reasoning tasks within this dataset typically consist of a descriptive scenario followed by a culminating question. Answering these questions requires performing multi-step mathematical calculations based on the context provided in the description.
D.3 Planning: ALFWorld
AlfWorld [39] is a comprehensive suite of synthetic, text-based environments set within a virtual household, featuring six distinct task types: Pick, Clean, Heat, Cool, Examine, and Pick Two. Each task presents a unique high-level objective (e.g., “put a vase in the safe”) that requires the agent to navigate and interact with various objects or receptacles (e.g., go to shelf 6, clean apple). To accomplish the assigned task, the agent must execute a series of actions to achieve the specified goal. However, the challenge lies in the object’s potential location - it could be in any of over 50 possible places within a given task instance - necessitating sequential exploration of each location by the agent. Consequently, the complete action sequence may encompass more than 50 discrete actions, posing a considerable challenge to the agent’s capabilities and efficiency.
Appendix E Baseline Details
E.1 LaMP: Personalization
We compare our proposed M-Pilot with several competitive baselines, encompassing both one-stage and two-stage methods. For all baseline approaches, we employ a consistent prompt template and utilize BM25 as the default retrieval mechanism across all experiments.
@@ E.1 LaMP: Personalization
For our ablation study, we primarily compare M-Pilot with the following ablated baseline:
  • M-Pilot w/o IDPO utilizes the controller model Llama-3-8B-Instruct to first generate a summary of the user’s retrieved history data. It then combines this summary with the input question as prompts for the environment model gpt-4o-mini to generate the final answer.
E.2 GSM: Reasoning
@@
  • M-Pilot w/o IDPO first utilizes a vanilla LLaMA3-8B-Instruct to break down the problem into sub-questions, and then gpt-3.5-turbo provide solutions based on both the main problem and the decomposed sub-questions.
@@
E.3 Alfworld: Planning
We compare M-Pilot with several strong baselines in the planning task, encompassing both one-stage and two-stage approaches. For all baselines, we employ gpt-3.5-turbo as the black-box model for task execution. The ablated baselines (two-stage) include w/o Guidance Optimization, w/o $1^{\text{st}},2^{\text{nd}}$-round IDPO, and w/o $2^{\text{nd}}$-round IDPO. Additional baselines (one-stage) include BUTLER [39], ReAct [55], Reflexion [38], and AdaPlanner [41].
@@ E.3 Alfworld: Planning
    • w/o $1^{\text{st}},2^{\text{nd}}$-round IDPO indicates a LLaMA3-8B-Instruct model that has undergone supervised fine-tuning on a limited amount of training data.
    • w/o $2^{\text{nd}}$-round IDPO denotes the LLaMA3-8B-Instruct model further trained using DPO on {positive, negative} pairs from the training set, building upon the supervised fine-tuned model.
  • BUTLER [39] is an agent that initially learns to perform abstract tasks in TextWorld through Imitation Learning (IL) and subsequently transfers the acquired policies to embodied tasks in ALFWorld.
  • ReAct [55] is a general paradigm that combines reasoning and acting with language models to solve diverse language reasoning and decision-making tasks.
  • Reflexion [38] employs verbal reinforcement to help agents learn from prior failures.
  • AdaPlanner [41] is a closed-loop planning method where the LLM plays two roles, planner and refiner. It leverages code-based prompting for precise planning and refinement.
Appendix F Implementation Details
F.1 Hardware and Software
@@
F.2 Training Cost
Table 19: Detailed API cost per 1 million tokens.
Backbone Model Input cost (\$) / 1M tokens Output cost (\$) / 1M tokens
gpt-3.5-turbo 3.0 6.0
gpt-4o-mini 0.15 0.6
The data generation cost was calculated by aggregating the total token consumption statistics provided by Azure API and subsequently applying the cost per token (gpt-3.5-turbo-0125, gpt-4o-mini) as specified in official documentation  [^3] https://openai.com/api/pricing/. The cost for processing 1M tokens, as detailed in Table 19, served as the basis for this calculation.
For the AlfWorld dataset, the entire training set consists of 8,808 samples. On average, using gpt-3.5-turbo to sample 100 examples costs approximately \$3.20, making the estimated cost for complete data collection \$282. For the GSM8K dataset, the full training set comprises 7,473 samples. The average cost for sampling 100 examples using gpt-3.5-turbo is \$1.215, resulting in an estimated total cost for data collection of \$90.80. In comparison, fine-tuning the gpt-3.5-turbo costs \$216.50, and requires hourly payment in deployment for inference. For LaMP-1, LaMP-2M, LaMP-2N, LaMP-3, LaMP-4, we use gpt-4o-mini for data generation. The total costs are separately \$6.144, \$1.882, \$2.348, \$8.111, \$10.022, with 5252, 2719, 2369, 8506, 12518 generated data samples.
During the training phase, we used four H100 GPUs for two rounds of DPO training. The process took approximately 1.5 hours for AlfWorld and GSM8K, resulting in a total training cost of 6 GPU hours. It took separately 4 gpu hours, 4 gpu hours, 4 gpu hours, 8 gpu hours, 12 gpu hours for the training process of LaMP-1, -2M, -2N, -3, -4.
F.3 LaMP: Personalization
F.3.1 Algorithm Details
We formalize the personalization problem within the context of our proposed M-Pilot framework. Specifically, we employ the controller model Llama-3-8B-Instruct to analyze the user’s retrieved history data and generate an informative and clear intermediate summary. This summary is then combined with the input question as prompts for the environment model gpt-4o-mini to derive the final answer. To enhance control capabilities, we utilize online DPO to optimize the controller model Llama-3-8B-Instruct.
@@ F.3.1 Algorithm Details
To prevent overfitting and reward hacking, the interaction stage processes the entire training dataset once for all personalization tasks. We sample at most two contrastive pairs for each training data point. We employ LoRA (Low-Rank Adaptation), a parameter-efficient method, to update the controller model Llama-3-8B-Instruct. LoRA is well-suited for personalization tasks, allowing efficient and effective optimization of the controller model. We utilize DPO for optimization.
F.3.2 Hyperparameter Configurations
We set the maximum sequence length for generated solutions to 512 tokens across all tasks and scenarios. The controller model is Llama-3-8B-Instruct, while the environment model is gpt-4o-mini for the primary tasks and gpt-3.5-turbo for specific ablation studies. For each user, we retrieve the minimum of $k$ and the total number of user profiles as historical profiles. The value of $k$ varies by dataset: all profiles for LaMP-1, 120 for LaMP-2N, 150 for LaMP-2M, 30 for LaMP-3, and 50 for LaMP-4. These retrieved profiles are utilized in generating intermediate solutions. Comprehensive prompt templates and additional details are provided in Appendix H.
To prevent overfitting and reward hacking, we iterate through the entire training dataset only once. For each data point, we perform ten interactions, generating ten distinct intermediate solutions with a temperature setting of 1.0. Consequently, each data point results in at least ten intermediate generations. To further mitigate overfitting and reward hacking, we sample a maximum of two contrastive pairs per data point. The total number of contrastive pairs sampled during the interaction stage is as follows: 5410 for LaMP 1, 2850 for LaMP-2M, 2548 for LaMP-2N, 4320 for LaMP-3, and 12518 for LaMP-4, respectively.
@@ F.3.2 Hyperparameter Configurations
F.4 GSM8K: Reasoning
Following the PAL framework [15], we employ code-style LLM prompts to facilitate the conversion of mathematical problems into executable code, thereby augmenting the model’s problem-solving capabilities. Unlike PAL, which directly translates mathematical problems into code, M-Pilot first assists GPT in decomposing the problem into more manageable sub-problems. This decomposition allows GPT to more effectively convert these simpler sub-problems into code, enhancing both the correctness and stability of the generated code. Additionally, since M-Pilot is responsible solely for high-level planning without engaging in low-level execution, we can train on the GSM8K dataset and evaluate on the GSM-Hard dataset. Both datasets comprise similar problem types, with GSM-Hard featuring more intricate numerical calculations.